# Transolver++: An Accurate Neural Solver for PDEs on Million-Scale Geometries

### v1

@@ 1 Introduction
Extensive physics processes can be precisely described by partial differential equations (PDEs) (Wazwaz2002PartialDE; evans2010partial), such as air dynamics of driving cars or internal stress of buildings. Accurately solving these PDEs is essential to industrial manufacturing (kopriva2009implementing; roubivcek2013nonlinear). However, it is hard and usually impossible to obtain the analytic solution of PDEs. Thus, numerical methods (solin2005partial) have been widely explored, whose typical solving process is first discretizing PDEs into computation meshes and then approximating solution on discretized meshes (grossmann2007numerical). In real applications, such as simulating a driving car or aircraft during takeoff, it will take several days or even months for calculation, and the simulation accuracy is highly affected by the fitness of discretized computation mesh (solanki2003finite; elrefaie2024drivaernet++). To speed up this process, deep models have been explored as efficient surrogates of numerical methods, known as neural PDE solvers (wang2023scientific). Training from pre-collected simulation data, neural solvers can learn to approximate the mapping between input and output of numerical methods and directly infer new samples in a flash (li2021fourier), posing a promising direction for industrial simulation.
Usually, industrial applications involve large and complex geometries, requiring the model to capture intricate physics processes underlying geometries efficiently. Although previous works have made some progress in handling complex geometries (li2023geometry; hao2023gnot) and attempted to speed up model efficiency, they are still far from real applications, especially in handling large geometries. Specifically, as illustrated in Figure 1(a), existing models fail to scale up beyond 400k points, while real applications typically involves million-scale mesh points or even more. Note that, as aforementioned, the fineness of computation meshes is the foundation of simulation accuracy. As the comparison of car mesh shown in Figure 1(b), limited mesh size will seriously sacrifice the precision of the geometry, where the originally streamlined surface becomes rough and uneven. This will further limit the simulation accuracy of physics processes, especially for aerodynamic applications (elrefaie2024drivaernet++). Thus, the capability of handling large geometries is indispensable for practical neural PDE solvers.
@@
As the latest progress in neural PDE solver architectures, Transolver (wu2024Transolver) proposes to learn intrinsic physical states underlying complex geometries and apply the attention mechanism among learned physical states to capture intricate physics interactions, which frees model capacities from unwieldy mesh points and achieves outstanding performance in car and airfoil simulations. Although Transolver demonstrates favorable capability in handling complex geometries, its maximum input size remains limited to 700k points (Figure 1(a)), and its experiment data is much simpler than real applications. Building on the essential step made by Transolver, we attempt to advance it to million-scale or even larger geometries in pursuing practical neural solvers.
@@ 2.1 Neural PDE Solver
Traditional numerical methods for solving PDEs often require high computational costs to achieve accurate solutions (solanki2003finite). Recently, deep learning methods have demonstrated remarkable potential as efficient surrogate models for solving PDEs due to their inherent non-linear modeling capability, known as neural PDE solvers.
@@
As a typical paradigm, operator learning has been widely studied for solving PDEs by learning the mapping between input functions and solutions. FNO (li2020neural) first proposes to approximate integral in the Fourier domain for PDE solving. Subsequently, Geo-FNO (li2021fourier) extends FNO to irregular meshes by transforming them into regular grids in the latent space. To further enhance the capabilities of FNO, U-FNO Wen2021UFNOA and U-NO rahman2022u are presented by leveraging U-Net (ronneberger2015u) to capture multiscale properties. Considering the high dimensionality of real-world PDEs, LSM (wu2023LSM) applies spectral methods in a learned lower-dimensional latent space to approximate input-output mappings. Afterward, LNO (wang2024LNO) adopts the attention mechanism to effectively map data from geometric space to latent space for complex geometries.
@@
Recently, Transformers (NIPS2017_3f5ee243) have achieved impressive progress in extensive fields and have also been applied to solving PDEs, where the attention mechanism has been proven as a Monte-Carlo approximation for global integral (jmlr_operator). However, standard attention suffers from quadratic complexity. Thus, many models like Galerkin (Cao2021ChooseAT), OFormer (li2023transformer) and FactFormer (anonymous2023factorized) propose different efficient attention mechanisms. Among them, GNOT (hao2023gnot) utilizes well-established linear attention, like Reformer or Performer (kitaev2020reformer; performer), and separately encodes geometric information, achieving favorable performance. However, linear attention often suffers from degraded performance as an approximation of standard attention (qin2022devil). Moreover, these Transformer-based methods treat input geometries as a sequence of mesh points and directly apply attention among mesh points, which may fall short in geometric learning and computation efficiency. As a significant advancement in PDE solving, Transolver (wu2024Transolver) introduces the Physics-Attention mechanism that groups massive mesh points into multiple physical states and applies attention among states, thereby enabling more effective and intrinsic modeling of complex physical correlations. However, it still faces challenges in degenerated physics learning and a high computation burden under million-scale geometries. These challenges will be well addressed in our paper.
@@
In addition to Transformers, graph neural networks (GNNs) (hamilton2017inductive; gao2019graph; pfaff2021learning) are also inherently suitable to process unstructured meshes by explicitly modeling the message passing among nodes and edges. GNO (Li2020NeuralOG) first implements neural operator with GNN. Later, GINO (li2023geometryinformed) combines GNO with Geo-FNO to encode the geometry information. Recently, 3D-GeoCA (anonymous2023geometryguided) integrates pre-trained 3D vision models to achieve a better representation learning of geometries. However, prone to geometric instability (instabilityGNN), GNNs can lead to unstable results and may be insufficient in capturing global physics interactions, especially when handling large-scale unstructured meshes (morris2023geometric).
@@ 2.2 Parallelism of Deep Models
Despite processing large-scale geometries being crucial in industrial design, this problem has not been explored in previous research. This paper breaks the computation bottleneck by leveraging the parallelism framework, which is related to the following seminal works. Towards general needs in handling the high GPU memory usage caused by large-scale inputs, several parallel frameworks have been proposed, such as tensor parallelism (shoeybi2019megatron) and model parallelism (huang2019gpipe). However, these methods are highly model-dependent and request significant communication overhead (zhuang2023optimizing). Another direction is to optimize attention mechanisms. Ring attention (liu2023ring), inspired by FlashAttention (dao2022flashattention), uses a ring topology between multi-GPUs, achieving quadratic communication complexity with respect to mesh points. Besides, DeepSpeed-Ulysses (jacobs2023deepspeed) splits the deep representation along the channel dimension and employs the All2All communication approach, reducing complexity to linear. Despite these improvements, the communication volume remains excessive. In contrast, Transolver++ leverages its unique physics-learning design and presents a highly optimized parallelism framework tailored to PDE solving, enabling minimal communication overhead and allowing for meshes of million-scale points.
@@ 3 Revisiting Transolver
Before Transolver++, we would like to briefly introduce Physics-Attention, the key design of Transolver (wu2024Transolver), and discuss its scaling issues for million-scale inputs.
@@
Finally, Physics-Attention employs the deslice operation to map updated states $\mathbf{s}^{\prime}$ back to mesh space using slice weights, i.e. $\mathbf{x}^{\prime}=\{\mathbf{x}_{i}^{\prime}\}_{i=1}^{N}=\{\sum_{j=1}^{M}\mathbf{w}_{ij}\mathbf{s}_{j}^{\prime}\}_{i=1}^{N}$. By replacing standard attention in Transformer (NIPS2017_3f5ee243) with Physics-Attention, we can obtain the Transolver.
@@ 4.1 Physics-Attention with Eidetic States
As aforementioned, if Physics-Attention is based on homogeneous physical states, it will degenerate to average pooling, which will damage the model’s performance. Thus, as the authors mentioned in Transolver (wu2024Transolver), they adopt the softmax function in calculating slice weights in Eq. (1), which can alleviate the indistinguishable states to some extent by guaranteeing peakier distribution (wu2022flowformer). However, we still observe the degeneration phenomenon when the model depth increases, as shown in Figure 2(b), which may result from the excessive focus on global information. Note that in industrial applications, the model’s capability of learning subtle physics is essential, which may highly affect the evaluation in industrial design (elrefaie2024drivaernet++), as a small diversion device can drastically change the wind drag of driving cars.
@@
As shown in Figure 2(b), modeling the distribution of each mesh point using a non-parametric approach has been proven to be impractical (ye2024differentialtransformer). Thus, we introduce a local adaptive mechanism that incorporates local information as a pointwise adjustment to the physics learning process. Specifically, we change the sharpness of the state distribution by learning to adjust the temperature $\tau_{0}$ in the Softmax function, namely
@@
As mentioned above, we try to learn eidetic physical states and assign mesh points to physical states with a possibility $\mathbf{w}$. In the canonical design of Transolver, it uses $\operatorname{Softmax}$ to form a categorical distribution across states. However, simply generating a categorical distribution is not enough, as we have not completely modeled the assignment process from points to certain physical states. Considering that direct sampling via $\operatorname{Argmax}$ is non-differentiable, we propose to adopt the Gumbel-Softmax (jang2017categorical) to perform differentiable sampling from the discrete categorical distribution, which is accomplished with the following reparameterization technique:
@@ 4.2 Parallel Transolver++
where $\oplus$ denotes the AllReduce operation (patarasuk2009bandwidth), which aggregates the results from all processes. In practice, the $k$-th GPU first independently computes its partial sums for the numerator $(\sum_{i=1}^{N_{k}}\mathbf{w}_{ij}^{(k)}\mathbf{x}_{i}^{(k)})$ and denominator $(\sum_{i=1}^{N_{k}}\mathbf{w}_{ij}^{(k)})$ of $N_{k}$ points. Next, these partial results are synchronized across GPUs to compute the eidetic states $\mathbf{s}_{j}$, highlighted by blue-marked steps in Algorithm 1.
@@ 5 Experiments
As summarized in Table 1, our experiments include both standard benchmarks and industrial simulations, which cover a broad range of mesh sizes. Specifically, the standard benchmarks include Elasticity, Plasticity, Airfoil, Pipe, NS2d, and Darcy, which are widely used in previous studies (wu2024Transolver). To further evaluate the model’s efficacy in real applications, we also perform experiments on industrial design tasks, where we utilized DrivAerNet++ (elrefaie2024drivaernet++) for car design and a newly simulated AirCraft dataset for 3D aircraft design. In addition to the error of predicted physics fields, we also measure the model performance for design by calculating drag and lift coefficients from predicted physics fields.
@@
We widely compare Transolver++ against more than 20 advanced baselines, covering various types of approaches. These baselines include 12 neural operators, such as Galerkin (Cao2021ChooseAT), LNO (wang2024LNO), and GINO (li2023geometryinformed), some of which are specifically designed for irregular meshes; 4 Transformer-based PDE solvers, including OFormer (li2023transformer), FactFormer (li2023scalable), GNOT (hao2023gnot), and Transolver (wu2024Transolver); and 4 graph-neural-network-based methods: GraphSAGE (hamilton2017inductive), PointNet (qi2017pointnet), Graph U-Net (gao2019graph), and MeshGraphNet (pfaff2021learning). Transolver is the previous state-of-the-art model. During experiments, we found that some neural operators designed for grids perform poorly for large-scale irregular meshes. Therefore, we only report their performance on standard benchmark datasets.
@@ Results
GraphSAGE (hamilton2017inductive) 0.328 0.284 0.282 0.859 0.294 0.040 0.988 0.109
PointNet (qi2017pointnet) 0.285 0.478 0.301 0.831 0.237 0.095 0.982 0.169
Graph U-Net (gao2019graph) 0.241 0.260 0.272 0.876 0.193 0.063 0.953 0.161
MeshGraphNet (pfaff2021learning) 0.529 0.422 0.260 0.870 0.209 0.038 0.993 0.113
GNO (li2020neural) 0.510 0.664 0.252 0.882 0.196 0.031 0.991 0.129
Galerkin (Cao2021ChooseAT) 0.234 0.274 0.267 0.792 0.235 0.069 0.879 0.118
Geo-FNO (Li2022FourierNO) 0.718 0.892 0.288 0.831 0.291 0.243 0.903 0.395
GINO (li2023geometryinformed) 0.586 0.638 0.323 0.725 0.220 0.047 0.983 0.133
GNOT (hao2023gnot) 0.174 0.171 0.158 0.901 0.167 0.033 0.991 0.093
LNO (wang2024LNO) 0.180 0.203 0.208 0.855 0.195 0.091 0.992 0.137
3D-GeoCA (anonymous2023geometryguided) 0.389 0.224 0.205 0.883 0.175 0.022 0.993 0.097
Transolver (wu2024Transolver) 0.173 0.167 0.061 0.931 0.145 0.037 0.994 0.092
@@ 5.2 PDEs on Large Geometries
To evaluate the performance in practical applications, we conduct experiments on two industrial datasets with million-scale meshes: DrivAerNet++ (elrefaie2024drivaernet++) and Aircraft. The latter is newly presented by us, which is of high quality and simulated by aerodynamicists. To fully evaluate the model performance under different scales, we split DrivAerNet++ into two scales, one only with surface pressure ($\sim$700k mesh points in each sample) and the other with full physics fields (2.5M mesh points).

### v2

@@ 1 Introduction
Extensive physics processes can be precisely described by partial differential equations (PDEs) (Wazwaz, 2002; Evans, 2010), such as air dynamics of driving cars or internal stress of buildings. Accurately solving these PDEs is essential to industrial manufacturing (Kopriva, 2009; Roubíček, 2013). However, it is hard and usually impossible to obtain the analytic solution of PDEs. Thus, numerical methods (Ŝolín, 2005) have been widely explored, whose typical solving process is first discretizing PDEs into computation meshes and then approximating solution on discretized meshes (Grossmann et al., 2007). In real applications, such as simulating a driving car or aircraft during takeoff, it will take several days or even months for calculation, and the simulation accuracy is highly affected by the fitness of discretized computation mesh (Solanki et al., 2003; Elrefaie et al., 2024). To speed up this process, deep models have been explored as efficient surrogates of numerical methods, known as neural PDE solvers (Wang et al., 2023). Training from pre-collected simulation data, neural solvers can learn to approximate the mapping between input and output of numerical methods and directly infer new samples in a flash (Li et al., 2021), posing a promising direction for industrial simulation.
Usually, industrial applications involve large and complex geometries, requiring the model to capture intricate physics processes underlying geometries efficiently. Although previous works have made some progress in handling complex geometries (Li et al., 2023b; Hao et al., 2023) and attempted to speed up model efficiency, they are still far from real applications, especially in handling large geometries. Specifically, as illustrated in Figure 1(a), existing models fail to scale up beyond 400k points, while real applications typically involves million-scale mesh points or even more. Note that, as aforementioned, the fineness of computation meshes is the foundation of simulation accuracy. As the comparison of car mesh shown in Figure 1(b), limited mesh size will seriously sacrifice the precision of the geometry, where the originally streamlined surface becomes rough and uneven. This will further limit the simulation accuracy of physics processes, especially for aerodynamic applications (Elrefaie et al., 2024). Thus, the capability of handling large geometries is indispensable for practical neural PDE solvers.
@@
As the latest progress in neural PDE solver architectures, Transolver (Wu et al., 2024) proposes to learn intrinsic physical states underlying complex geometries and apply the attention mechanism among learned physical states to capture intricate physics interactions, which frees model capacities from unwieldy mesh points and achieves outstanding performance in car and airfoil simulations. Although Transolver demonstrates favorable capability in handling complex geometries, its maximum input size remains limited to 700k points (Figure 1(a)), and its experiment data is much simpler than real applications. Building on the essential step made by Transolver, we attempt to advance it to million-scale or even larger geometries in pursuing practical neural solvers.
@@ 2.1 Neural PDE Solver
Traditional numerical methods for solving PDEs often require high computational costs to achieve accurate solutions (Solanki et al., 2003). Recently, deep learning methods have demonstrated remarkable potential as efficient surrogate models for solving PDEs due to their inherent non-linear modeling capability, known as neural PDE solvers.
@@
As a typical paradigm, operator learning has been widely studied for solving PDEs by learning the mapping between input functions and solutions. FNO (Li et al., 2020a) first proposes to approximate integral in the Fourier domain for PDE solving. Subsequently, Geo-FNO (Li et al., 2021) extends FNO to irregular meshes by transforming them into regular grids in the latent space. To further enhance the capabilities of FNO, U-FNO Wen et al. (2022) and U-NO Rahman et al. (2023) are presented by leveraging U-Net (Ronneberger et al., 2015) to capture multiscale properties. Considering the high dimensionality of real-world PDEs, LSM (Wu et al., 2023) applies spectral methods in a learned lower-dimensional latent space to approximate input-output mappings. Afterward, LNO (Wang & Wang, 2024) adopts the attention mechanism to effectively map data from geometric space to latent space for complex geometries.
@@
Recently, Transformers (Vaswani et al., 2017) have achieved impressive progress in extensive fields and have also been applied to solving PDEs, where the attention mechanism has been proven as a Monte-Carlo approximation for global integral (Kovachki et al., 2023). However, standard attention suffers from quadratic complexity. Thus, many models like Galerkin (Cao, 2021), OFormer (Li et al., 2023c) and FactFormer (Tran et al., 2023) propose different efficient attention mechanisms. Among them, GNOT (Hao et al., 2023) utilizes well-established linear attention, like Reformer or Performer (Kitaev et al., 2020; Choromanski et al., 2021), and separately encodes geometric information, achieving favorable performance. However, linear attention often suffers from degraded performance as an approximation of standard attention (Qin et al., 2022). Moreover, these Transformer-based methods treat input geometries as a sequence of mesh points and directly apply attention among mesh points, which may fall short in geometric learning and computation efficiency. As a significant advancement in PDE solving, Transolver (Wu et al., 2024) introduces the Physics-Attention mechanism that groups massive mesh points into multiple physical states and applies attention among states, thereby enabling more effective and intrinsic modeling of complex physical correlations. However, it still faces challenges in degenerated physics learning and a high computation burden under million-scale geometries. These challenges will be well addressed in our paper.
@@
In addition to Transformers, graph neural networks (GNNs) (Hamilton et al., 2017; Gao & Ji, 2019; Pfaff et al., 2021) are also inherently suitable to process unstructured meshes by explicitly modeling the message passing among nodes and edges. GNO (Li et al., 2020b) first implements neural operator with GNN. Later, GINO (Li et al., 2023a) combines GNO with Geo-FNO to encode the geometry information. Recently, 3D-GeoCA (Deng et al., 2024) integrates pre-trained 3D vision models to achieve a better representation learning of geometries. However, prone to geometric instability (Klabunde & Lemmerich, 2023), GNNs can lead to unstable results and may be insufficient in capturing global physics interactions, especially when handling large-scale unstructured meshes (Morris et al., 2023).
@@ 2.2 Parallelism of Deep Models
Despite processing large-scale geometries being crucial in industrial design, this problem has not been explored in previous research. This paper breaks the computation bottleneck by leveraging the parallelism framework, which is related to the following seminal works. Towards general needs in handling the high GPU memory usage caused by large-scale inputs, several parallel frameworks have been proposed, such as tensor parallelism (Shoeybi et al., 2019) and model parallelism (Huang et al., 2019). However, these methods are highly model-dependent and request significant communication overhead (Zhuang et al., 2023). Another direction is to optimize attention mechanisms. Ring attention (Liu et al., 2023), inspired by FlashAttention (Dao et al., 2022), uses a ring topology between multi-GPUs, achieving quadratic communication complexity with respect to mesh points. Besides, DeepSpeed-Ulysses (Jacobs et al., 2023) splits the deep representation along the channel dimension and employs the All2All communication approach, reducing complexity to linear. Despite these improvements, the communication volume remains excessive. In contrast, Transolver++ leverages its unique physics-learning design and presents a highly optimized parallelism framework tailored to PDE solving, enabling minimal communication overhead and allowing for meshes of million-scale points.
@@ 3 Revisiting Transolver
Before Transolver++, we would like to briefly introduce Physics-Attention, the key design of Transolver (Wu et al., 2024), and discuss its scaling issues for million-scale inputs.
@@
Finally, Physics-Attention employs the deslice operation to map updated states $\mathbf{s}^{\prime}$ back to mesh space using slice weights, i.e. $\mathbf{x}^{\prime}=\{\mathbf{x}_{i}^{\prime}\}_{i=1}^{N}=\{\sum_{j=1}^{M}\mathbf{w}_{ij}\mathbf{s}_{j}^{\prime}\}_{i=1}^{N}$. By replacing standard attention in Transformer (Vaswani et al., 2017) with Physics-Attention, we can obtain the Transolver.
@@ 4.1 Physics-Attention with Eidetic States
As aforementioned, if Physics-Attention is based on homogeneous physical states, it will degenerate to average pooling, which will damage the model’s performance. Thus, as the authors mentioned in Transolver (Wu et al., 2024), they adopt the softmax function in calculating slice weights in Eq. (1), which can alleviate the indistinguishable states to some extent by guaranteeing peakier distribution (Wu et al., 2022). However, we still observe the degeneration phenomenon when the model depth increases, as shown in Figure 2(b), which may result from the excessive focus on global information. Note that in industrial applications, the model’s capability of learning subtle physics is essential, which may highly affect the evaluation in industrial design (Elrefaie et al., 2024), as a small diversion device can drastically change the wind drag of driving cars.
@@
As shown in Figure 2(b), modeling the distribution of each mesh point using a non-parametric approach has been proven to be impractical (Ye et al., 2024). Thus, we introduce a local adaptive mechanism that incorporates local information as a pointwise adjustment to the physics learning process. Specifically, we change the sharpness of the state distribution by learning to adjust the temperature $\tau_{0}$ in the Softmax function, namely
@@
As mentioned above, we try to learn eidetic physical states and assign mesh points to physical states with a possibility $\mathbf{w}$. In the canonical design of Transolver, it uses $\operatorname{Softmax}$ to form a categorical distribution across states. However, simply generating a categorical distribution is not enough, as we have not completely modeled the assignment process from points to certain physical states. Considering that direct sampling via $\operatorname{Argmax}$ is non-differentiable, we propose to adopt the Gumbel-Softmax (Jang et al., 2017) to perform differentiable sampling from the discrete categorical distribution, which is accomplished with the following reparameterization technique:
@@ 4.2 Parallel Transolver++
where $\oplus$ denotes the AllReduce operation (Patarasuk & Yuan, 2009), which aggregates the results from all processes. In practice, the $k$-th GPU first independently computes its partial sums for the numerator $(\sum_{i=1}^{N_{k}}\mathbf{w}_{ij}^{(k)}\mathbf{x}_{i}^{(k)})$ and denominator $(\sum_{i=1}^{N_{k}}\mathbf{w}_{ij}^{(k)})$ of $N_{k}$ points. Next, these partial results are synchronized across GPUs to compute the eidetic states $\mathbf{s}_{j}$, highlighted by blue-marked steps in Algorithm 1.
@@ 5 Experiments
As summarized in Table 1, our experiments include both standard benchmarks and industrial simulations, which cover a broad range of mesh sizes. Specifically, the standard benchmarks include Elasticity, Plasticity, Airfoil, Pipe, NS2d, and Darcy, which are widely used in previous studies (Wu et al., 2024). To further evaluate the model’s efficacy in real applications, we also perform experiments on industrial design tasks, where we utilized DrivAerNet++ (Elrefaie et al., 2024) for car design and a newly simulated AirCraft dataset for 3D aircraft design. In addition to the error of predicted physics fields, we also measure the model performance for design by calculating drag and lift coefficients from predicted physics fields.
@@
We widely compare Transolver++ against more than 20 advanced baselines, covering various types of approaches. These baselines include 12 neural operators, such as Galerkin (2021), LNO (2024), and GINO (2023a), some of which are specifically designed for irregular meshes; 4 Transformer-based PDE solvers, including OFormer (2023c), FactFormer (2023d), GNOT (2023), and Transolver (2024); and 4 graph-neural-network-based methods: GraphSAGE (2017), PointNet (2017), Graph U-Net (2019), and MeshGraphNet (2021). Transolver is the previous state-of-the-art model. During experiments, we found that some neural operators designed for grids perform poorly for large-scale irregular meshes. Therefore, we only report their performance on standard benchmark datasets.
@@ Results
GraphSAGE (2017) 0.328 0.284 0.282 0.859 0.294 0.040 0.988 0.109
PointNet (2017) 0.285 0.478 0.301 0.831 0.237 0.095 0.982 0.169
Graph U-Net (2019) 0.241 0.260 0.272 0.876 0.193 0.063 0.953 0.161
MeshGraphNet (2021) 0.529 0.422 0.260 0.870 0.209 0.038 0.993 0.113
GNO (2020a) 0.510 0.664 0.252 0.882 0.196 0.031 0.991 0.129
Galerkin (2021) 0.234 0.274 0.267 0.792 0.235 0.069 0.879 0.118
Geo-FNO (2022) 0.718 0.892 0.288 0.831 0.291 0.243 0.903 0.395
GINO (2023a) 0.586 0.638 0.323 0.725 0.220 0.047 0.983 0.133
GNOT (2023) 0.174 0.171 0.158 0.901 0.167 0.033 0.991 0.093
LNO (2024) 0.180 0.203 0.208 0.855 0.195 0.091 0.992 0.137
3D-GeoCA (2024) 0.389 0.224 0.205 0.883 0.175 0.022 0.993 0.097
Transolver (2024) 0.173 0.167 0.061 0.931 0.145 0.037 0.994 0.092
@@ 5.2 PDEs on Large Geometries
To evaluate the performance in practical applications, we conduct experiments on two industrial datasets with million-scale meshes: DrivAerNet++ (Elrefaie et al., 2024) and Aircraft. The latter is newly presented by us, which is of high quality and simulated by aerodynamicists. To fully evaluate the model performance under different scales, we split DrivAerNet++ into two scales, one only with surface pressure ($\sim$700k mesh points in each sample) and the other with full physics fields (2.5M mesh points).